How do I translate this into equations to be solved? Thanks!
A bricklayer and an electrician together spend 90 hours working on a new house. If the bricklayer earns $12 per hour, the electrician earns $16 an hour, and the owner pays them a total of $1350 for their work, how many hours does each worker spend on the house?

Answer by snovember85(4)

Let x = hours that the bricklayer worked, y = hours that the electrician worked
know: (1) x + y = 90 and (2) 12x + 16y = 1350
x + y = 90 => x = 90 - y
Replace x by 90 - y in equation (2), so 12(90 - y) + 16y = 1350
Solve this for y. then plug it back in equation (1) to get x.